20.50 1800 N e F 2.2 10 Specity the 2.0 eC 5.0 cm 180 :5.0cm 50cm -20aC 20 C
Answer #1

Q.20.24)

Enet = 2*E*cos(45 deg)

NOTE: The vertical components of the electric fields will cancel out and the horizontal components will add up

So, Enet = 2*(k*q/r^2)*cos(45 deg)

= 2*(9*10^9*1*10^-9/(0.05^2 + 0.05^2))*cos(45 deg)

= 2545.6 N/C

b)

direction = 0 deg (the net electric field will be horizontal)

2)

Enet = 4*(k*q/r^2)*cos(45 deg)

= 4* (9*10^9*2*10^-9/(0.02^2 + 0.02^2))*cos(45 deg)

= 63639 N/C

So, force = Enet*q = 63639*(3.5*10^-9)

= 2.23*10^-4 N <----------- your answer's unit is wrong i guess. It should be in N